the length of a rectangle is 8 in. more then the width. the perimeter of the rectangle is 24 in. what is the width and length of the rectangle.
Answer by rfer(16322)

2x+2(x+8)=24
2x+2x+16=24
4x=8
x=2in
x+8=10in
2(2)+2(2+8)=24
4+4+16=24